Calculate the order of magnitude ?( ) of the following functions, assuming that T (1) = ? (1).

T(n) = 5T(n/7) + nlogn

T(n) = 3T(n/4) + log3n

T(n) = 4T(n/4) + nlogn

T(n) = 9T(n/10) + log3n

T(n) = 2T(n/3) + n/log2n

T(n) = T(n-1) + 1/n
Answer #1

Answer:

1. T(n) = 5T(n/7) + nlogn

Using Master's theorm , we have :

a = 5 , b = 7 , c = 1 , k = n

log a base b = log 5 base 7 = 0.827

log a base b < c

Therefore T(n) = theta(n^c) = theta(n^0.827) = theta(n)

2. T(n) = 3T(n/4) + log3n

Using Master's theorm , we have :

a = 3 , b = 4 , c = 0 , k = 3

log a base b = log 3 base 4 = 0.792

log a base b > c

therefore T(n) = theta(f(n)) = theta(log3n)

3.  T(n) = 4T(n/4) + nlogn

Here a = 4 , b = 4 , c = 1 , k = 1

log4 base 4 = 1 , here log a base b = c

Therefore , T(n) = theta(n^clog^k+1n) = theta(n^1log^k+1n) = theta(nlog^2n

4.  T(n) = 9T(n/10) + log3n

Here a = 9 , b = 10 , c = 0 , k = 3

log 9 base 10 = 1 , c = 0

here loga base b > c , therefore T(n) = theta(f(n) = theta(log3n)